# The Unbounded Denominators Conjecture

### v3

@@ 4.5. Invariant vectors
We recall the following definition (cf. [CV19, §3.7]).
@@
If $\widehat{\Gamma}$ denotes the congruence completion of the group $\Gamma$, then $H^{1,\mathrm{cong}}(\Gamma,\mathbf{F}_{\ell})\simeq H^{1}(\widehat{\Gamma},\mathbf{F}_{\ell})$. For a prime $\ell$, one may define ([CE12]) the groups
$\widetilde{H}^{1}(\mathbf{F}_{\ell}):=\lim_{N}H^{1}(\Gamma(N),\mathbf{F}_{\ell}),\quad\widetilde{H}^{1}(\mathbf{Q}/\mathbf{Z}):=\lim_{N}H^{1}(\Gamma(N),\mathbf{Q}/\mathbf{Z}).$
@@
Let $N$ be an integer,and $\eta\in H^{1}(\Gamma(N),\mathbf{F}_{\ell})$. If, for all $g\in\mathrm{SL}_{2}(\mathbf{Z}/N\mathbf{Z})$, the class $g\eta-\eta\in H^{1}(\Gamma(N),\mathbf{F}_{\ell})$ is a congruence, then $\eta$ is congruence.
@@
It suffices to prove the vanishing of $H_{2}(\Delta,\mathbf{Z})$ for any Sylow subgroup of $\mathrm{SL}_{2}(\mathbf{F}_{p})$. For odd primes, the Sylow subgroup is cyclic and the cohomology of a cyclic group is only non-zero in even degree. The $2$-Sylow subgroup is a generalized quaternion group, whose cohomology also vanishes in odd degree (as follows from [Hup67, Satz 25.3(a), p.643] and [Swa60, Theorem 2]). ∎
@@
and thus the result follows from Lemma 4.5.5. Thus we may assume that $\ell=p$. Assume that $p>2$. Let $G(p)$ be the $p$-congruence subgroup of $\mathrm{SL}_{2}(\mathbf{Z}_{p})$. We may assume that $p>2$. Then $G(p)$ is $p$-torsion free and $p$-powerful, so, with $M=M_{0}(\mathbf{F}_{p})^{\vee}$ where $G(p)/G(p^{2})\simeq M_{0}(\mathbf{F}_{p})$, we deduce by Lazard’s Theorem ([Laz65, Chapter V, 2.2.6.3 and 2.2.7.2, page 167]) that there are isomorphisms
@@
where the cup product map $\wedge^{2}M:H^{1}\wedge H^{1}\rightarrow H^{2}$ is an isomorphism. Assuming $p\geq 3$, we find that $M\simeq M^{\vee}$ is self-dual as a $\mathrm{SL}_{2}(\mathbf{F}_{p})$-module and so $\wedge^{2}M\simeq M$. Moreover, we have  $M^{\mathrm{SL}_{2}(\mathbf{F}_{p})}=0$. Consider the Hochschild–Serre spectral sequence:
$H^{i}(\mathrm{SL}_{2}(\mathbf{F}_{p}),H^{j}(G(p),\mathbf{F}_{p}))\Rightarrow H^{i+j}(\mathrm{SL}_{2}(\mathbf{Z}_{p}),\mathbf{F}_{p}).$
Since $H^{1}(\mathrm{SL}_{2}(\mathbf{F}_{p}),\mathbf{F}_{p})=H^{1}(\mathrm{SL}_{2}(\mathbf{Z}_{p}),\mathbf{F}_{p})$ and $M^{\mathrm{SL}_{2}(\mathbf{F}_{p})}=0$, there is an exact sequence:
@@
If $\ell$ is odd and $N$ is a non-trivial power of $\ell$ or $N\geq 8$ is a power of $\ell=2$, then the map
@@
Here $H^{2}(\Gamma(N),\mathbf{F}_{\ell})=0$ because $\Gamma(N)$ is a free group. The last vertical map is zero by the previous lemma if $N\in\{3,5,8,16\}$, and thus the image of $(\widetilde{H}^{1})^{\mathrm{SL}_{2}(\widehat{\mathbf{Z}})}$ in $(\widetilde{H}^{1})^{\widehat{\Gamma}(N)}$ lands in the image of $H^{1}(\widehat{\Gamma}(N),\mathbf{F}_{\ell})$ in these cases. But these are finite groups we can compute explicitly.
@@
$\dim H^{1}(\Gamma(4),\mathbf{F}_{2})^{\mathrm{SL}_{2}(\mathbf{Z}/4\mathbf{Z})}=\dim H^{1}(\Gamma(8),\mathbf{F}_{2})^{\mathrm{SL}_{2}(\mathbf{Z}/8\mathbf{Z})}=\dim H^{1}(\Gamma(16),\mathbf{F}_{2})^{\mathrm{SL}_{2}(\mathbf{Z}/16\mathbf{Z})}=1.$
@@
Recall that for $N=3$, 44, and $5$ that $X(N)$ has genus zero. Hence the cohomology $V=H^{1}(\Gamma(N),\mathbf{Z})$ is coming entirely from the the cusps, which correspond to the cosets of $\langle\left(\begin{matrix}1&1\\ 0&1\end{matrix}\right)\rangle$ in $\mathrm{PSL}_{2}(\mathbf{Z}/N\mathbf{Z})$. In particular, in the Grothendieck group,
@@
    For $N=4$, we have $\mathrm{PSL}_{2}(\mathbf{Z}/4\mathbf{Z})=S_{4}$, and $V_{\mathbf{Q}}$ is a sum of absolutely irreducible representations of dimensions $2$ and $3$. The group $S_{4}$ has two Brauer characters of dimension $1$ and $2$ respectively. The $2$ dimensional representation remains irreducible and the semi-simplification of both the $3$ dimensional representations has constituents of dimensions $1$ and $2$. Hence the dimension of the invariant space of $V^{\mathrm{ss}}_{\mathbf{F}_{\ell}}$ is $1$-dimensional. But $H^{1}(\Gamma(4)/\Gamma(8),\mathbf{F}_{2})$ is a direct sum of the $1$ and $2$ dimensional representations, so this one-dimensional constituent occurs as a sub-representation.
@@
Let $v\in\widetilde{H}^{1}(\mathbf{F}_{\ell})^{\mathrm{SL}_{2}(\widehat{\mathbf{Z}})}$. From the Künneth formula and Lemma 4.5.6, we deduce that there is an isomorphism $H^{2}(\mathrm{SL}_{2}(\widehat{\mathbf{Z}}),\mathbf{F}_{\ell})=0$ for any prime $\ell>5$. Thus, for $\ell>5$, any class $v$ lies in the image of $H^{1}(\mathrm{SL}_{2}(\mathbf{Z}),\mathbf{F}_{\ell})\simeq\mathbf{Z}/12\mathbf{Z}\otimes\mathbf{F}_{\ell}$, but one easily checks that the map
@@
is an isomorphism, and thus $v=0$.
Suppose that $\ell=3$ or $\ell=5$. Since the map $H^{2}(\mathrm{SL}_{2}(\widehat{\mathbf{Z}}),\mathbf{F}_{\ell})\rightarrow H^{2}(\widehat{\Gamma}(\ell),\mathbf{F}_{\ell})$ is zero by Lemma 4.5.9, the image of $v$ lands in the $\mathrm{SL}_{2}(\mathbf{Z}/\ell\mathbf{Z})$-invariants of $H^{1}(\Gamma(\ell),\mathbf{F}_{\ell})$. Thus it contributes to the $\mathrm{SL}_{2}(\mathbf{Z}/\ell\mathbf{Z})$-invariants of the semi-simplification of $H^{1}(\Gamma(\ell),\mathbf{F}_{\ell})$ as a $\mathrm{SL}_{2}(\mathbf{Z}/\ell\mathbf{Z})$-module. But this space has dimension zero by Lemma 4.5.10.
@@
is zero, and thus the element $v$ lies in the image of $H^{1}(\Gamma(8),\mathbf{F}_{2})$. Furthermore, it has the property that the $\mathrm{SL}_{2}(\mathbf{Z}/8\mathbf{Z})$-module generated by $v$ is trivial after passing to the quotient by the congruence homology
$H^{1}(\widehat{\Gamma}(8),\mathbf{F}_{2})\simeq H^{1}(\widehat{\Gamma}(8)/\widehat{\Gamma}(16),\mathbf{F}_{2}).$
But that means that the image of $v$ in $H^{1}(\widehat{\Gamma}(16),\mathbf{F}_{2})$ is genuinely invariant under $\mathrm{SL}_{2}(\mathbf{Z}/16\mathbf{Z})$. By Lemma 4.5.10, the space of such invariants is one dimensional. But this one dimensional space lands in the image of $H^{1}(\widehat{\Gamma}(16),\mathbf{F}_{2})$, and thus $v$ is trivial in $\widetilde{H}^{1}(\mathbf{F}_{2})$. ∎
@@
From Theorem 4.5.6, it suffices to consider the case of $n=2$ and $\ell=p\in\{2,3,5\}$. There is an exact sequence:
$0\rightarrow H_{2}(\mathrm{SL}_{2}(\mathbf{Z}_{p}),\mathbf{Z})/\ell\rightarrow H_{2}(\mathrm{SL}_{2}(\mathbf{Z}_{p}),\mathbf{F}_{\ell})\rightarrow H_{1}(\mathrm{SL}_{2}(\mathbf{Z}_{p}),\mathbf{Z})[\ell]\rightarrow 0.$
Since $H_{1}(\mathrm{SL}_{2}(\mathbf{Z}_{p}),\mathbf{Z})\simeq\mathbf{Z}/12\mathbf{Z}\otimes\mathbf{Z}_{p}$, this implies the desired lower bounds for in Lemma 4.5.11, with equality in all cases if and only if $H_{2}(\mathrm{SL}_{2}(\mathbf{Z}_{p}),\mathbf{Z})/\ell=0$ for all primes $\ell$. Since the group $H_{2}(\mathrm{SL}_{2}(\mathbf{Z}_{p}),\mathbf{Z})$ is a finitely generated abelian group, to deduce that $H_{2}(\mathrm{SL}_{2}(\mathbf{Z}_{p}),\mathbf{Z})=0$ it suffices to prove the upper bounds.
For the upper bounds, we deduce from Theorem 4.5.2 that there is an injection
$H^{2}(\mathrm{SL}_{2}(\mathbf{Z}_{p}),\mathbf{F}_{p})\rightarrow H^{2}(\mathrm{SL}_{2}(\widehat{\mathbf{Z}}),\mathbf{F}_{p})\rightarrow H^{2}(\mathrm{SL}_{2}(\mathbf{Z}),\mathbf{F}_{p}).$
But for $n>0$ we have
@@
from which it follows that $H^{2}(\mathrm{SL}_{2}(\mathbf{Z}),\mathbf{F}_{p})=\mathbf{F}_{p}$ if $p\in\{2,3\}$ and is zero otherwise. This completes the proof. ∎

### v4

@@ 4.5. Invariant vectors
The congruence completion $\widehat{\Gamma}$ of a congruence subgroup $\Gamma\subset\mathrm{SL}_{2}(\mathbf{Z})$ is the inverse limit of all quotients of $\Gamma$ by normal congruence subgroups. We recall the following definition (cf. [CV19, § 3.7]).
@@
If $\widehat{\Gamma}$ denotes the congruence completion of the group $\Gamma$, then $H^{1,\mathrm{cong}}(\Gamma,\mathbf{F}_{\ell})\simeq H^{1}(\widehat{\Gamma},\mathbf{F}_{\ell})$. In practice, we shall usually talk about $H^{1}(\widehat{\Gamma},\mathbf{F}_{\ell})$ rather than $H^{1,\mathrm{cong}}(\Gamma,\mathbf{F}_{\ell})$ but we have recalled the definition here to allow for an easier comparison with the arguments of [CV19]. For a prime $\ell$, one may define ([CE11, § 2], [CE16, § 1], see also [CE12]) the groups
$\widetilde{H}^{1}(\mathbf{F}_{\ell}):=\lim_{N}H^{1}(\Gamma(N),\mathbf{F}_{\ell}),\quad\widetilde{H}^{1}(\mathbf{Q}/\mathbf{Z}):=\lim_{N}H^{1}(\Gamma(N),\mathbf{Q}/\mathbf{Z})$
@@
Let $N$ be an integer, and $\eta\in H^{1}(\Gamma(N),\mathbf{F}_{\ell})$. If, for all $g\in\mathrm{SL}_{2}(\mathbf{Z}/N\mathbf{Z})$, the class $g\eta-\eta\in H^{1}(\Gamma(N),\mathbf{F}_{\ell})$ is a congruence class, then $\eta$ is a congruence class.
@@
It suffices to prove the vanishing of $H_{2}(\Delta,\mathbf{Z})$ for any Sylow subgroup $\Delta$ of $\mathrm{SL}_{2}(\mathbf{F}_{p})$. For odd primes, the Sylow subgroup is cyclic and the cohomology of a cyclic group is only non-zero in even degree. For a finite group $G$, we have $H^{n+1}(G,\mathbf{Z})\simeq\mathrm{Ext}^{1}(H_{n}(G,\mathbf{Z}),\mathbf{Z})$ for $n\geq 0$ by the universal coefficient theorem. Hence the homology of a cyclic group is zero in even degree $n>0$. The $2$-Sylow subgroup is a generalized quaternion group, whose cohomology also vanishes in odd degree (as follows from [Hup67, Satz 25.3(a), p.643] and [Swa60, Theorem 2]), and once more we are done by the universal coefficient theorem. ∎
@@
and thus the result follows from Lemma 4.5.5. Thus we may assume that $\ell=p$. Assume that $p>2$. Let $G(p)$ be the $p$-congruence subgroup of $\mathrm{SL}_{2}(\mathbf{Z}_{p})$. Recall that a group $G$ is $p$-powerful if $[G,G]$ is contained in the subgroup generated by $p$th powers (for $p$ odd) or $4$th powers for $p=2$. The group $G(p)$ is $p$-torsion free and $p$-powerful (for $p>2$), so, with $M=M_{0}(\mathbf{F}_{p})^{\vee}$ where $G(p)/G(p^{2})\simeq M_{0}(\mathbf{F}_{p})$, we deduce by Lazard’s Theorem ([Laz65, Chapter V, 2.2.6.3 and 2.2.7.2, page 551]) that there are isomorphisms
@@
where the cup product map $\wedge^{2}M:H^{1}\wedge H^{1}\rightarrow H^{2}$ is an isomorphism. Assuming $p\geq 3$, we find that $M\simeq M^{\vee}$ is self-dual as a $\mathrm{SL}_{2}(\mathbf{F}_{p})$-module and so $\wedge^{2}M\simeq M$. Moreover, we have an equality $M^{\mathrm{SL}_{2}(\mathbf{F}_{p})}=0$. Consider the Hochschild–Serre spectral sequence:
$E^{2}_{i,j}=H^{i}(\mathrm{SL}_{2}(\mathbf{F}_{p}),H^{j}(G(p),\mathbf{F}_{p}))\Rightarrow H^{i+j}(\mathrm{SL}_{2}(\mathbf{Z}_{p}),\mathbf{F}_{p}).$
Since $M^{\mathrm{SL}_{2}(\mathbf{F}_{p})}=0$ and $M\simeq\wedge^{2}M$ we have $E^{2}_{0,1}=E^{2}_{0,2}=0$. It follows that $E^{\infty}_{0,1}=E^{\infty}_{0,2}=0$, but also that $E^{\infty}_{2,0}=E^{2}_{2,0}=H^{2}(\mathrm{SL}_{2}(\mathbf{F}_{p}),\mathbf{F}_{p})$. The vanishing of $E^{\infty}_{0,2}$ implies that $H^{2}(\mathrm{SL}_{2}(\mathbf{Z}_{p}),\mathbf{F}_{p})$ is an extension of $E^{\infty}_{2,0}$ by $E^{\infty}_{1,1}\subseteq E^{2}_{1,1}=H^{1}(\mathrm{SL}_{2}(\mathbf{F}_{p}),M)$, and hence there is an exact sequence:
@@
If $\ell$ is odd and $N$ is a nontrivial power of $\ell$ or $N\geq 8$ is a power of $\ell=2$, then the map
@@
Here $H^{2}(\Gamma(N),\mathbf{F}_{\ell})=0$ because $\Gamma(N)$ is a free group. (For $N\geq 3$, $\Gamma(N)$ is torsion free, and so $\Gamma(N)$ may be identified with the fundamental group of a surface $\mathbf{H}/\Gamma(N)$ with cusps.) The last vertical map is zero by the previous lemma if $N\in\{3,5,8,16\}$, and thus the image of $(\widetilde{H}^{1})^{\mathrm{SL}_{2}(\widehat{\mathbf{Z}})}$ in $(\widetilde{H}^{1})^{\widehat{\Gamma}(N)}$ lands in the image of $H^{1}(\widehat{\Gamma}(N),\mathbf{F}_{\ell})$ in these cases. But these are finite groups we can compute explicitly.
@@
$\dim H^{1}(\Gamma(4),\mathbf{F}_{2})^{\mathrm{SL}_{2}(\mathbf{Z}/4\mathbf{Z})}=$ $1.\displaystyle\dim H^{1}(\Gamma(8),\mathbf{F}_{2})^{\mathrm{SL}_{2}(\mathbf{Z}/8\mathbf{Z})}=\dim H^{1}(\Gamma(6),\mathbf{F}_{2})^{\mathrm{SL}_{2}(\mathbf{Z}/16\mathbf{Z})}=$1.
@@
Recall that for $N=3$, 44, and $5$ that $X(N)$ has genus zero. Hence the cohomology of the module $V=H^{1}(\Gamma(N),\mathbf{Z})$ is coming entirely from the the cusps, which correspond to the cosets of $\langle\left(\begin{matrix}1&1\\ 0&1\end{matrix}\right)\rangle$ in $\mathrm{PSL}_{2}(\mathbf{Z}/N\mathbf{Z})$. In particular, in the Grothendieck group $K_{0}(\mathbf{Q}[\mathrm{SL}_{2}(\mathbf{Z}/N\mathbf{Z})])$ of $\mathrm{SL}_{2}(\mathbf{Z}/N\mathbf{Z})$-representations over $\mathbf{Q}$,
@@
    For $N=4$, we have $\mathrm{PSL}_{2}(\mathbf{Z}/4\mathbf{Z})=S_{4}$, and $V_{\mathbf{Q}}$ is a sum of absolutely irreducible representations of dimensions $2$ and $3$. The group $S_{4}$ has two Brauer characters of dimension $1$ and $2$ respectively. The $2$-dimensional representation remains irreducible and the semi-simplification of both the $3$-dimensional representations has constituents of dimensions $1$ and $2$. Hence the invariant space of $V^{\mathrm{ss}}_{\mathbf{F}_{\ell}}$ is $1$-dimensional. But $H^{1}(\Gamma(4)/\Gamma(8),\mathbf{F}_{2})$ is a direct sum of the $1$ and $2$-dimensional representations, so this $1$-dimensional constituent occurs as a sub-representation.
@@
We need to show that any $v\in\widetilde{H}^{1}(\mathbf{F}_{\ell})^{\mathrm{SL}_{2}(\widehat{\mathbf{Z}})}$ is zero. Let us consider the images of $v$ under various maps in equation (4.5.10). We first note that $H^{1}(\mathrm{SL}_{2}(\mathbf{Z}),\mathbf{F}_{\ell})\simeq\mathbf{Z}/12\mathbf{Z}\otimes\mathbf{F}_{\ell}$ and that the map
@@
is an isomorphism for any $\ell$. Hence we may assume the image of $v$ in $H^{2}(\mathrm{SL}_{2}(\widehat{\mathbf{Z}}),\mathbf{F}_{\ell})$ is non-zero. From the Künneth formula and Lemma 4.5.6, there is an isomorphism $H^{2}(\mathrm{SL}_{2}(\widehat{\mathbf{Z}}),\mathbf{F}_{\ell})=0$ for any prime $\ell>5$, and hence we may assume that $\ell\leq 5$.
Suppose that $\ell=3$ or $\ell=5$, and take $N=\ell$ in equation (4.5.10). We proved that the map $H^{2}(\mathrm{SL}_{2}(\widehat{\mathbf{Z}}),\mathbf{F}_{\ell})\rightarrow H^{2}(\widehat{\Gamma}(\ell),\mathbf{F}_{\ell})$ is zero by Lemma 4.5.9. It follows that the image of $v$ in $(\widetilde{H}^{1}(\mathbf{F}_{\ell}))^{\widehat{\Gamma}(\ell)}$ is $\mathrm{SL}_{2}(\mathbf{Z}/\ell\mathbf{Z})$-invariant and lands in the image of $H^{1}(\Gamma(\ell),\mathbf{F}_{\ell})$. Thus the $\mathrm{SL}_{2}(\mathbf{Z}/\ell\mathbf{Z})$-invariants of the semi-simplification of $H^{1}(\Gamma(\ell),\mathbf{F}_{\ell})$ as a $\mathrm{SL}_{2}(\mathbf{Z}/\ell\mathbf{Z})$-module is nontrivial. But this space has dimension 0 by Lemma 4.5.11.
@@
is zero, and thus, arguing as in the case $\ell=3$ or $5$ above, the image of $v$ in $(\widetilde{H}^{1}(\mathbf{F}_{2}))^{\widehat{\Gamma}(8)}$ coincides with the image of an element $w\in H^{1}(\Gamma(8),\mathbf{F}_{2})$. Furthermore, the $\mathrm{SL}_{2}(\mathbf{Z}/8\mathbf{Z})$-module generated by $w$ is $\mathrm{SL}_{2}(\mathbf{Z}/8\mathbf{Z})$-invariant after passing to the quotient by the congruence homology
$H^{1}(\widehat{\Gamma}(8),\mathbf{F}_{2})\simeq H^{1}(\widehat{\Gamma}(8)/\widehat{\Gamma}(16),\mathbf{F}_{2})\simeq H^{1}(\Gamma(8)/\Gamma(16),\mathbf{F}_{2}).$
But that means that the image of $w$ in $H^{1}(\Gamma(16),\mathbf{F}_{2})$ is invariant under $\mathrm{SL}_{2}(\mathbf{Z}/16\mathbf{Z})$. By Lemma 4.5.11, the space of such invariants is $1$-dimensional. But this $1$-dimensional space lands in the image of $H^{1}(\widehat{\Gamma}(16),\mathbf{F}_{2})$, and thus the image of $w$ and hence also of $v$ must be trivial in $\widetilde{H}^{1}(\mathbf{F}_{2})^{\widehat{\Gamma}(16)}\subset\widetilde{H}^{1}(\mathbf{F}_{2})$, and in particular $v=0$. ∎
@@
From Lemma 4.5.6, it suffices to consider the case of $n=2$ and $\ell=p$. For any $\ell$ and $p$, there is an exact sequence:
(4.5.13) $0\rightarrow H_{2}(\mathrm{SL}_{2}(\mathbf{Z}_{p}),\mathbf{Z})/\ell\rightarrow H_{2}(\mathrm{SL}_{2}(\mathbf{Z}_{p}),\mathbf{F}_{\ell})\rightarrow H_{1}(\mathrm{SL}_{2}(\mathbf{Z}_{p}),\mathbf{Z})[\ell]\rightarrow 0.$
Since $H_{1}(\mathrm{SL}_{2}(\mathbf{Z}_{p}),\mathbf{Z})\simeq\mathbf{Z}/12\mathbf{Z}\otimes\mathbf{Z}_{p}$, this proves that $H^{2}(\mathrm{SL}_{2}(\mathbf{Z}_{p}),\mathbf{F}_{p})$ has dimension at least one when $p\in\{2,3\}$. Suppose we prove that $H^{2}(\mathrm{SL}_{2}(\mathbf{Z}_{p}),\mathbf{F}_{\ell})$ has dimension at most one when $\ell=p\in\{2,3\}$ and dimension zero otherwise. First, this would complete the computation of $H^{2}(\mathrm{SL}_{2}(\mathbf{Z}_{p}),\mathbf{F}_{\ell})$. Second, it would follow that the second map in equation (4.5.13) is always an isomorphism, and so $H_{2}(\mathrm{SL}_{2}(\mathbf{Z}_{p}),\mathbf{Z})/\ell=0$ for all primes $\ell$. Since the group $H_{2}(\mathrm{SL}_{2}(\mathbf{Z}_{p}),\mathbf{Z})$ is a finitely generated abelian group, this will also show that that $H_{2}(\mathrm{SL}_{2}(\mathbf{Z}_{p}),\mathbf{Z})=0$, completing the proof of the lemma.
Let us now bound from above the dimension of $H^{2}(\mathrm{SL}_{2}(\mathbf{Z}_{p}),\mathbf{F}_{\ell})$. We may assume that $\ell=p$ by Lemma 4.5.6. There are maps:
(4.5.14) $H^{2}(\mathrm{SL}_{2}(\mathbf{Z}_{\ell}),\mathbf{F}_{\ell})\simeq H^{2}(\mathrm{SL}_{2}(\widehat{\mathbf{Z}}),\mathbf{F}_{{\ell}})\hookrightarrow H^{2}(\mathrm{SL}_{2}(\mathbf{Z}),\mathbf{F}_{{\ell}}),$
where the first map is an isomorphism by Lemma (4.5.9), and the second map is an inclusion from the exact sequence (4.5.10) and the vanishing of $(\widetilde{H}^{1}(\mathbf{F}_{\ell}))^{\mathrm{SL}_{2}(\widehat{\mathbf{Z}})}$ by Theorem 4.5.2. But for $n>0$ we have
@@
from which it follows that $H^{2}(\mathrm{SL}_{2}(\mathbf{Z}),\mathbf{F}_{\ell})=\mathbf{F}_{\ell}$ if $\ell\in\{2,3\}$ and is zero otherwise. This gives the desired upper bound on the dimension of $H^{2}(\mathrm{SL}_{2}(\mathbf{Z}_{\ell}),\mathbf{F}_{\ell})$ via the inclusion (4.5.14), completing the proof. ∎